OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (8,-4) ; slope: -5/6

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{5}{6} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values }\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

How do i find m ?

OpenStudy (jdoe0001):

\(\large \bf slope = {\color{green}{ m}}= -\cfrac{5}{6}\)

OpenStudy (jdoe0001):

is given

OpenStudy (anonymous):

but that seems way too easy

OpenStudy (jdoe0001):

it's

OpenStudy (anonymous):

whoo thanks <33

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

oh but wait , the onyl answer where both y-y1 and ( x-x1) are subtraction is the one with 5/6 instead of -5/6 D. y-4=5/6(x-8)